OpenStudy (gabylovesyou):

The equation of line EF is y = 2x + 1. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0,2) @ganeshie8 @johnweldon1993

OpenStudy (btaylor):

In order for a line to be parallel it must have the same slope. Therefore, the slope will be 2. So we have y=2x+b Given the point, plug in those x and y values and you'll get B, the y-intercept.

OpenStudy (gabylovesyou):

2 = 2(0) + b ?

OpenStudy (compassionate):

"The equation of line EF is y = 2x + 1. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0,2) " Okay so use the formula y - y1 = m(x - x1) Parallel lines have equal slopes. So m (slope) will be 2 in this, too. So plug in your x-y points (0 , 2) y - y1 = m(x - x1) y - 2 = 2(x - 0) [because our slope is parallel, so 2 is m, y1 is our y-axis, so that's 2, and x1 is our x-axis, so that's 0.) Now... y - 2 = 2(x - 0) [[Distribute]] y - 2 = 2x [2 * 0 just equals 0, so we leave it out and just have 2x.] Now add 2 to both sides to solve for y. y - 2 + 2 = 2x + 2 y = 2x + 2 And wala! There you go :)